Evaluate the integral of (sin x + cos x)² dx


1
Expert's answer
2021-03-15T04:58:47-0400

Evaluate the integral of (sin x + cos x)² dx

Solution. Simplify the integral function:

"( sinx+cosx)^2=sin^2x+cos^2x+2sinxcosx=1+sin(2x)."

Then: "\intop(sinx+cosx)^2dx=\intop(1+sin(2x))dx=x-\frac{1}{2}cos(2x)+\rm constant."

Answer: "\intop(sinx+cosx)^2dx=x-\frac{1}{2}cos(2x)+\rm constant."